OpenStudy (anonymous):

What percent of 81 is 506.25 ?

sammixboo (sammixboo):

\(\Large \rm \color{#20bd23}{\dfrac{part}{whole}=\dfrac{percent}{100}}\) Part = 506.25 Whole = 81 Percent = Unknown Plug in the values \(\Large \rm \color{#20bd23}{\dfrac{506.25}{81}=\dfrac{percent}{100}}\) Cross multiply 506.25 by 100, so what is 506.25 x 100?